# Flavor solitons in dense neutrino gases

### v1

@@ I Introduction
[…]
@@
[…]
@@ II.2 Pendulum Solutions
An instability requires a spectral crossing, i.e., $G_{v}$ must change sign for some $-1<v_{\rm c}<+1$. At present, we are only examining spectra that have a single crossing. This necessary condition would also be sufficient if we were to include inhomogeneous normal modes, i.e., those with a non-vanishing wave number Morinaga:2021vmc . For the homogeneous case, a sufficient condition based on the Nyquist criterion is (Appendix A)
[…]
@@
[…]
@@
[…]
@@ II.3 Homogeneous Temporal Soliton
[…]
@@
[…]
@@ II.5 Uniformly Moving Soliton
We further redefine $v^{\prime}=k^{\prime}/\omega^{\prime}=(v-V)/(1-vV)$ as the velocity in the new frame. The range $-1\leq v\leq 1$ of course maps on $-1\leq v^{\prime}\leq 1$ as behooves a Lorentz transformation, and conversely $v=(v^{\prime}+V)/(1+v^{\prime}V)$. Moreover, we redefine the Bloch vectors in the new frame
@@
In the laboratory frame, the solution depends then only on the combination $t^{\prime}=\gamma(t-Vr)$, or, equivalently on the combination $r-t/V$. We recognize this as a uniformly moving wavefront with the velocity $v_{\mathrm{soliton}}=1/V$. Since $|V|<1$, it follows that the corresponding soliton is superluminal. The homogeneous temporal soliton is obtained from here in the limit $V=0$, which formally corresponds to a soliton moving with infinite speed.
@@ II.7 Dispersion Relation
For our discrete three-mode examples, one may simply evaluate the dispersion relation of Eq. (19) directly. In analogy to the lab-frame case, it is a quadratic equation for $\Omega^{\prime}$. For our example A, one finds a non-vanishing imaginary part for $-0.099<V<0.40$, for B the range is $-0.43<V<0.43$. Even though the two extreme neutrino beams have $v=\pm 1$, the allowed range for $V$ does not exhaust this full range.
@@ III.1 Static Soliton
which, except for a change in sign, has the same form as Eq. (6), an observation first made by Johns et al. Johns:2019izj . Compared to the temporal case, in which the variable $v$ runs in the compact interval from $-1$ to $1$, in this EOM the role of the variable $v$ is played by $v^{-1}$, which lies in the non-compact interval $|v^{-1}|>1$. However, this modification does not lead to significant differences in the physics and the nature of the solutions.
@@ III.2 Uniformly Moving Soliton
This is now identical to Eq. (22), with the variable $v^{\prime}$ still defined in the range between $-1$ and $1$. Therefore, it admits solutions which behave as a static soliton in the boosted reference frame. Such solutions depend only on the variable $r^{\prime}=\gamma(r-Vt)$. Therefore, in the laboratory frame, they are solutions depending both on space and time, corresponding to a soliton moving with the subluminal velocity $V$.
@@ A.1 Nyquist Criterion for Homogeneous Case
[…]
As we have shown in Ref. Fiorillo:2023mze , this dispersion relation admits the same unstable frequencies. In other words the zeros of $\phi(\Omega)$ and $\tilde{\phi}(\Omega)$ in the upper half of the complex plane defined by Im$(\Omega)>0$ are the same.
@@
[…]
@@
Therefore, as $\Omega$ grows from $\infty$- ∞ to +$\infty$+ ∞, the function $\phi(\Omega)$ starts from 00 and returns to 00 after having changed its sign and wrapped itself around the origin $N$ times.
We first confirm the validity of the criterion empirically in two simple cases, the angular distributions for cases A and D of Ref. Padilla-Gay:2021haz . For both cases, we adjust the sign of the distribution such that $G_{1}>0$. Figure 9 shows the trajectories drawn by $\phi(\Omega)$ in the complex plane as $\Omega$ runs from $\infty$- ∞ to +$\infty$+ ∞. At $\Omega\to-\infty$, $\phi(\Omega)\to G_{1}/\Omega$, so that $\phi$ starts from 00 and moves along the negative real axis. At $\Omega\to+\infty$, $\phi(\Omega)$ returns again to the origin from the positive real axis, therefore confirming our result in Eq. (41) that it acquires a phase of $\pi$.
@@
Figure 9: Nyquist diagrams for the angular distribution of case A (left) and case D (right) of Ref. Padilla-Gay:2021haz . We show the trajectory of the function $\phi(\Omega)$ in the complex plane, as $\Omega$ evolves from $\infty$- ∞ to +$\infty$+ ∞ along the real axis. Since the angular distribution has non-analytical discontinuities at $v=-1$ and $v=+1$, we smooth them out introducing a rapid cutoff at these positions.
We can now derive a simple criterion to determine whether a single-crossed spectrum $G_{v}$ admits an instability without explicitly solving the dispersion relation. The key idea is that the function $\phi(\Omega)$ can only cross the real axis, for real $\Omega$, either at $\Omega=0$ or at $\Omega=-v_{\rm c}G_{1}$, where $v_{\rm c}$ is the position of the crossing $G_{v_{\rm c}}=0$. Without loss of generality, we choose $G_{1}>0$. Therefore, at $\Omega\to-\infty$, the function $\phi(\Omega)\to G_{1}/\Omega$ starts from 00 tangent to the negative real axis. At $\Omega\to+\infty$, $\phi(\Omega)$ returns to 00 tangent to the positive real axis. The number of times that it wraps around the origin in between corresponds to the number of unstable frequencies. In order to wrap around the origin once, $\phi(\Omega)$ must cross the real axis first on the positive side, as in both panels of Fig. 9. Afterwards, to surround the origin, it should cross it again on the negative side, as in the right panel of Fig. 9, so that it can finally return to the positive side and asymptotically reach zero. It is easy to see that with one crossing, $\phi(\Omega)$ can wrap around the origin at most once, so there is at most one unstable frequency.
@@ A.2 Superluminal Soliton
[…]
@@ B.1 Equations of Motion
We briefly review the gyroscopic pendulum, following Appendix B of Ref. Raffelt:2011yb . This contraption, also known as symmetric heavy top, Lagrangian top, or spherical pendulum with spin, is an axially symmetric body, spinning around its symmetry axis (moment of inertia $I_{s}$) with support on this axis Klein-Sommerfeld:2008 . Its moment of inertia relative to that point is $I$, mass $M$, gravitational acceleration $g$ along the $z$–direction, distance $\ell$ between support and center of mass, and angle $\vartheta$ relative to the $z$–direction, i.e., this angle is counted relative to the upward direction. The potential energy is $V=Mg\,\ell\,\cos\vartheta$. If the top were essentially point like, one would have $I=M\ell^{2}$, but in general one considers an extended body, where the center of mass and the “center of oscillation” are different.
@@
The total angular momentum ${\bf J}={\bf L}+{\bf S}$ has conserved $z$–component, where $S_{z}=S\,\cos\vartheta$ and $L_{z}=I\dot{\varphi}\sin^{2}\vartheta$. Here one factor of $\sin\vartheta$ comes from the projection of ${\bf r}$ on the transverse plane and the velocity is $\dot{\varphi}\sin\vartheta$. Therefore, $J_{z}=I\dot{\varphi}\sin^{2}\vartheta+S\cos\vartheta$ is conserved and
@@ B.2 Soliton Solution
In the context of the flavor pendulum, we are interested in motions that consist of a precession and a nutation between two limiting zenith angles $0\leq\vartheta_{1}\leq\vartheta\leq\vartheta_{2}\leq\pi$, corresponding to $11\geq c_{1}\geq c\geq c_{2}\geq-$ 1. In particular, if the highest position is $c_{1}=1$, the upright orientation, then $J_{z}=S$ and $j_{z}=1$ as well as $\varepsilon=1$. Moreover, we may absorb $\lambda$ in the definition of time, implying
@@
which indeed varies between $-1\leq c\leq+1$ for $0\leq\sigma\leq 1$. If this lowest point occurs at $t=0$, the solution is
@@
which varies as $-1<\tau<+1$ for $-\infty<t<\infty$.
@@
The complete orbital motion of the gyroscope is encoded in the polar angles given in Eq. (B.2). We may write the solution for the unit vector providing the orientation of the gyroscope in the form
@@
The derivatives can now be obtained by explicit differentiation of the solutions in Eq. (B.2).
@@ C.1 From Three Beams to Pendulum and Soliton
as stated in Eq. (II.2) of the main text. This eigenfrequency represents two complex conjugate solutions (and thus a true “soliton frequency”) only if the argument of the square root is positive, which also implies that the first term under the square root must be positive. In this case follows the identification of the pendulum natural frequency $\lambda=\sqrt{v_{1}v_{2}v_{3}G_{0}G_{1}}$ and spin parameter $\sigma=S/2\lambda$ stated in Eq. (12) of the main text.
@@ C.2 From Soliton to Three Beams
One immediately identifies $S=J_{z}=2\omega_{\rm P}=2\,{\rm Re}\,\Omega$. Moreover, from Eq. (II.2) we recall that $\lambda^{2}=\omega_{\rm P}^{2}+\Gamma^{2}=v_{1}v_{2}v_{3}G_{0}G_{1}$ so that overall the connections are
@@
Therefore, the remaining five three-mode parameters are very degenerate. We may pick $-1\leq v_{1}<v_{2}<v_{3}\leq+1$ and $G_{0}$ and $G_{1}$ anyway we like such that their product provides the desired $|\Omega|^{2}$.
@@ C.3 Explicit Solution for the Entire Spectrum
where $w=uG_{1}$ is the precession frequency of ${\bf L}_{u}$ around ${\bf D}_{1}$. The asymptotic state is ${\bf r}=\hat{\bf z}$ and ${\bf J}=S\,\hat{\bf z}$ and so the conserved length is $|{\bf L}_{u}|=|\lambda^{2}+w^{2}+wS|$ that we could use to normalize it to unity.
@@
where $c(t)$ was provided in Eq. (B.2) in terms of the pendulum natural frequency $\lambda$ and spin parameter $\sigma$. The asymptotic value is $c(\pm\infty)=1$ and $D_{v}(\pm\infty)=G_{v}$. By construction, the lowest soliton point occurs at $t=0$ and is $c(0)=2\sigma^{2}-1$. Therefore, the largest excursion of ${\bf D}_{v}$ from its asymptotic state is

### v2

@@ I Introduction
Another exact solution of the nonlinear flavor wave equation in the form of a uniformly moving flavor wave was recently discovered by Duan, Martin and Omanakuttan Duan:2021woc . We believe that this solution corresponds to the homogeneous pure-precession solution found in Ref. Raffelt:2007cb as seen from a boosted frame in the same way as our superluminal soliton corresponds to the homogeneous flavor pendulum as seen from a boosted frame.
@@
$\left(\partial_{t}+v\partial_{r}\right)\,{\bf D}_{v}=\mu\int_{-1}^{+1}dv^{\prime}\,\left(1-vv^{\prime}\right){\bf D}_{v^{\prime}}\times{\bf D}_{v},$ (1)
where a dependence on $(r,t)$ is implied for ${\bf D}_{v}$. The scale $\mu=\sqrt{2}G_{\rm F}(n_{\nu}+n_{\bar{\nu}})$ is a measure of the neutrino-neutrino refractive effect. The equivalent EOM in matrix form is $i(\partial_{t}+v\partial_{r})\,{\sf D}_{v}=\mu\int dv^{\prime}\,(1-vv^{\prime})[{\sf D}_{v^{\prime}},{\sf D}_{v}]$.
@@
$i\left(\partial_{t}+v\partial_{r}\right){\cal D}_{v}=\mu\int_{-1}^{+1}dv^{\prime}\,G_{v^{\prime}}\left(1-vv^{\prime}\right)\left({\cal D}_{v}-{\cal D}_{v^{\prime}}\right).$ (3)
@@ II.2 Pendulum Solutions
An instability requires a spectral crossing, i.e., $G_{v}$ must change sign for some $1<v_{\rm c}<+1$. At present, we are only examining spectra that have a single crossing. This necessary condition would also be sufficient if we were to include inhomogeneous normal modes, i.e., those with a non-vanishing wave number Morinaga:2021vmc . For the homogeneous case, a sufficient condition based on the Nyquist criterion is (Appendix A)
$\frac{G_{1}}{G_{0}v_{\rm c}}<0\quad\hbox{and}\quad\int dv\,\frac{vG_{v}}{G_{0}(v-v_{\rm c})}<0.$ (10)
@@
Figure 2: Pendulum trajectories in the $D^{x}_{1}$$D^{y}_{1}$ plane for the two benchmark cases defined in Table 1. Shown are two zenith-angle periods. The azimuth-angle period is generally not commensurate, causing the trajectory to fill the plane after many zenith periods.
@@
On the other hand, the pendulum trajectory is not periodic in general, as one can see in Fig. 2, where we show the projection of the pendulum trajectory in the $D^{x}_{1}$$D^{y}_{1}$ plane for two periods of $D_{1}^{z}$. The motion is actually conditionally periodic: $\vartheta(t)$ and $\varphi(t)$ are separately periodic, however with non-commensurate periods unless the seed was fine-tuned. We still refer to the pendular motion as periodic, with the implication that it is only periodic in the zenith and azimuth angle separately. The azimuth motion $\varphi(t)$ does not vanish even in the upright position except for vanishing spin (plane pendulum). The motion $\varphi(t)$ is never retrograde (Appendix B), the overall direction being determined by the initial spin orientation.
@@ II.3 Homogeneous Temporal Soliton
Figure 4: Soliton trajectories in the $D^{x}_{1}$$D^{y}_{1}$ plane for the two benchmark cases defined in Table 1.
@@
We may regard the soliton as the elementary object which is periodically repeated in the solutions of Figs. 2 and 2. For a soliton, we show in Figs. 4 and 4 the components $D^{z}_{1}$, $|D^{xy}_{1}|$, and $\dot{\varphi}$, as well as the trajectories projected in the $D^{x}_{1}$$D^{y}_{1}$ plane.
@@ II.5 Uniformly Moving Soliton
We further redefine $v^{\prime}=k^{\prime}/\omega^{\prime}=(v-V)/(1-vV)$ as the velocity in the new frame. The range $1\leq v\leq 1$ of course maps on $1\leq v^{\prime}\leq 1$ as behooves a Lorentz transformation, and conversely $v=(v^{\prime}+V)/(1+v^{\prime}V)$. Moreover, we redefine the Bloch vectors in the new frame
@@
In the laboratory frame, the solution depends then only on the combination $t^{\prime}=\gamma(t-Vr)$, or, equivalently on the combination rt/Vr-t/V. We recognize this as a uniformly moving wavefront with the velocity $v_{\mathrm{soliton}}=1/V$. Since $|V|<1$, it follows that the corresponding soliton is superluminal. The homogeneous temporal soliton is obtained from here in the limit $V=0$, which formally corresponds to a soliton moving with infinite speed.
@@ II.7 Dispersion Relation
For our discrete three-mode examples, one may simply evaluate the dispersion relation of Eq. (19) directly. In analogy to the lab-frame case, it is a quadratic equation for $\Omega^{\prime}$. For our example A, one finds a non-vanishing imaginary part for 0.099<V<0.40-0.099<V<0.40, for B the range is 0.43<V<0.43-0.43<V<0.43. Even though the two extreme neutrino beams have $v=\pm 1$, the allowed range for $V$ does not exhaust this full range.
@@ III.1 Static Soliton
which, except for a change in sign, has the same form as Eq. (6), an observation first made by Johns et al. Johns:2019izj . Compared to the temporal case, in which the variable $v$ runs in the compact interval from 1-1 to $1$, in this EOM the role of the variable $v$ is played by $v^{-1}$, which lies in the non-compact interval $|v^{-1}|>1$. However, this modification does not lead to significant differences in the physics and the nature of the solutions.
@@ III.2 Uniformly Moving Soliton
This is now identical to Eq. (22), with the variable $v^{\prime}$ still defined in the range between 1-1 and $1$. Therefore, it admits solutions which behave as a static soliton in the boosted reference frame. Such solutions depend only on the variable $r^{\prime}=\gamma(r-Vt)$. Therefore, in the laboratory frame, they are solutions depending both on space and time, corresponding to a soliton moving with the subluminal velocity $V$.
@@ A.1 Nyquist Criterion for Homogeneous Case
$\tilde{\phi}(\Omega)=\int\frac{vG_{v}}{vG_{1}+\Omega}\,dv=0,$ (36)
where $G_{1}=\int dv\,vG_{v}$ as in the main text. In our previous paper Fiorillo:2023mze , we have discussed how to deal with the pole $v=-\Omega/G_{1}$ when $\Omega$ is real. Since we are interested in unstable frequencies only, we choose to define the function on the real axis as
$\phi(\Omega)$ $=$ $\int\frac{vG_{v}}{vG_{1}+\Omega+i\epsilon}\,dv$ (37)
$=$ $\int_{\mathrm{PV}}\frac{vG_{v}}{vG_{1}+\Omega}\,dv+i\frac{\pi\Omega}{G_{1}^{2}}G_{-\Omega/G_{1}}.$
@@
$I=\int_{-\infty}^{+\infty}\frac{d\log\phi(\Omega)}{d\Omega}\,d\Omega$ (38)
@@
Therefore, as $\Omega$ grows from -$\infty$ to $\infty$, the function $\phi(\Omega)$ starts from 0 and returns to 0 after having changed its sign and wrapped itself around the origin $N$ times.
We first confirm the validity of the criterion empirically in two simple cases, the angular distributions for cases A and D of Ref. Padilla-Gay:2021haz . For both cases, we adjust the sign of the distribution such that $G_{1}>0$. Figure 9 shows the trajectories drawn by $\phi(\Omega)$ in the complex plane as $\Omega$ runs from -$\infty$ to $\infty$. At $\Omega\to$-$\infty$, $\phi(\Omega)\to G_{1}/\Omega$, so that $\phi$ starts from 0 and moves along the negative real axis. At $\Omega\to+\infty$, $\phi(\Omega)$ returns again to the origin from the positive real axis, therefore confirming our result in Eq. (41) that it acquires a phase of $\pi$.
@@
Figure 9: Nyquist diagrams for the angular distribution of case A (left) and case D (right) of Ref. Padilla-Gay:2021haz . We show the trajectory of the function $\phi(\Omega)$ in the complex plane, as $\Omega$ evolves from -$\infty$ to $\infty$ along the real axis. Since the angular distribution has non-analytical discontinuities at v=1v=-1 and $v=+1$, we smooth them out introducing a rapid cutoff at these positions.
We can now derive a simple criterion to determine whether a single-crossed spectrum $G_{v}$ admits an instability without explicitly solving the dispersion relation. The key idea is that the function $\phi(\Omega)$ can only cross the real axis, for real $\Omega$, either at $\Omega=0$ or at $\Omega=-v_{\rm c}G_{1}$, where $v_{\rm c}$ is the position of the crossing $G_{v_{\rm c}}=0$. Without loss of generality, we choose $G_{1}>0$. Therefore, at $\Omega\to$-$\infty$, the function $\phi(\Omega)\to G_{1}/\Omega$ starts from 0 tangent to the negative real axis. At $\Omega\to+\infty$, $\phi(\Omega)$ returns to 0 tangent to the positive real axis. The number of times that it wraps around the origin in between corresponds to the number of unstable frequencies. In order to wrap around the origin once, $\phi(\Omega)$ must cross the real axis first on the positive side, as in both panels of Fig. 9. Afterwards, to surround the origin, it should cross it again on the negative side, as in the right panel of Fig. 9, so that it can finally return to the positive side and asymptotically reach zero. It is easy to see that with one crossing, $\phi(\Omega)$ can wrap around the origin at most once, so there is at most one unstable frequency.
@@ A.2 Superluminal Soliton
$\phi(\Omega^{\prime})=\int\frac{v^{\prime}S^{z,0}_{v^{\prime}}}{v^{\prime}G^{\prime}_{1}+\Omega^{\prime}}d\xi^{\prime}=0.$ (42)
@@ B.1 Equations of Motion
We briefly review the gyroscopic pendulum, following Appendix B of Ref. Raffelt:2011yb . This contraption, also known as symmetric heavy top, Lagrangian top, or spherical pendulum with spin, is an axially symmetric body, spinning around its symmetry axis (moment of inertia $I_{s}$) with support on this axis Klein-Sommerfeld:2008 . Its moment of inertia relative to that point is $I$, mass $M$, gravitational acceleration $g$ along the zz–direction, distance $\ell$ between support and center of mass, and angle $\vartheta$ relative to the zz–direction, i.e., this angle is counted relative to the upward direction. The potential energy is $V=Mg\,\ell\,\cos\vartheta$. If the top were essentially point like, one would have $I=M\ell^{2}$, but in general one considers an extended body, where the center of mass and the “center of oscillation” are different.
@@
The total angular momentum ${\bf J}={\bf L}+{\bf S}$ has conserved zz–component, where $S_{z}=S\,\cos\vartheta$ and $L_{z}=I\dot{\varphi}\sin^{2}\vartheta$. Here one factor of $\sin\vartheta$ comes from the projection of $\bf r$} on the transverse plane and the velocity is $\dot{\varphi}\sin\vartheta$. Therefore, $J_{z}=I\dot{\varphi}\sin^{2}\vartheta+S\cos\vartheta$ is conserved and
@@ B.2 Soliton Solution
In the context of the flavor pendulum, we are interested in motions that consist of a precession and a nutation between two limiting zenith angles $0\leq\vartheta_{1}\leq\vartheta\leq\vartheta_{2}\leq\pi$, corresponding to $1\geq c_{1}\geq c\geq c_{2}\geq-1$. In particular, if the highest position is $c_{1}=1$, the upright orientation, then $J_{z}=S$ and $j_{z}=1$ as well as $\varepsilon=1$. Moreover, we may absorb $\lambda$ in the definition of time, implying
@@
which indeed varies between $1\leq c\leq+1$ for $0\leq\sigma\leq 1$. If this lowest point occurs at $t=0$, the solution is
@@
which varies as $1<\tau<+1$ for $\infty<t<\infty$.
@@
The complete orbital motion of the gyroscope is encoded in the polar angles given in Eq. (63). We may write the solution for the unit vector providing the orientation of the gyroscope in the form
@@
The derivatives can now be obtained by explicit differentiation of the solutions in Eq. (63).
@@ C.1 From Three Beams to Pendulum and Soliton
as stated in Eq. (11) of the main text. This eigenfrequency represents two complex conjugate solutions (and thus a true “soliton frequency”) only if the argument of the square root is positive, which also implies that the first term under the square root must be positive. In this case follows the identification of the pendulum natural frequency $\lambda=\sqrt{v_{1}v_{2}v_{3}G_{0}G_{1}}$ and spin parameter $\sigma=S/2\lambda$ stated in Eq. (12) of the main text.
@@ C.2 From Soliton to Three Beams
One immediately identifies $S=J_{z}=2\omega_{\rm P}=2\,{\rm Re}\,\Omega$. Moreover, from Eq. (13) we recall that $\lambda^{2}=\omega_{\rm P}^{2}+\Gamma^{2}=v_{1}v_{2}v_{3}G_{0}G_{1}$ so that overall the connections are
@@
Therefore, the remaining five three-mode parameters are very degenerate. We may pick $1\leq v_{1}<v_{2}<v_{3}\leq+1$ and $G_{0}$ and $G_{1}$ anyway we like such that their product provides the desired $|\Omega|^{2}$.
@@ C.3 Explicit Solution for the Entire Spectrum
where $w=uG_{1}$ is the precession frequency of ${\bf L}_{u}$ around ${\bf D}_{1}$. The asymptotic state is $^{\bf r}=\hat{\bf z}$ and $^{\bf J}=S\,\hat{\bf z}$ and so the conserved length is $|{\bf L}_{u}|=|\lambda^{2}+w^{2}+wS|$ that we could use to normalize it to unity.
@@
where $c(t)$ was provided in Eq. (63) in terms of the pendulum natural frequency $\lambda$ and spin parameter $\sigma$. The asymptotic value is $c(\pm\infty)=1$ and $D_{v}(\pm\infty)=G_{v}$. By construction, the lowest soliton point occurs at $t=0$ and is $c(0)=2\sigma^{2}-1$. Therefore, the largest excursion of ${\bf D}_{v}$ from its asymptotic state is